​ if 4/a = a^2/16, then check whether a is rational or irrational number.

if 4/a = a^2/16, then check whether a is rational or irrational number.

Grade:9

1 Answers

Debarati
2 years ago
According to the question, 4/a= a2/16
=> a3= 16*4= 4*4*4
=> 3\sqrt[]{}a33\sqrt[]{}4*4*4
=> a= 4
Since 4 is rational, a is a rational number.

I hope you understood.